OpenStudy (anonymous):

solve this equation pls and show work if you can 4k^2=5k ^= to the 2nd power

OpenStudy (anonymous):

4k²=5k 4k²-5k=0 k(4k-5)=0 k=0 4k-5=0 4k=5 k=5/4